• Derive an approximate formula for the speed of a satellite in a circular orbit at an altitude of 500 km from the surface of the Earth. How long does it take for the satellite to complete one orbit?
1
Expert's answer
2020-06-09T13:21:01-0400

According to Newton's law of gravitation, satellite moves around the earth due to centripetal force by earth on satellite.

Let, speed of satellite is "v" ,mass is "m" and height from surface of the earth is "h=500km=5\times10^{5}m"


Thus,

"G\frac{M_em}{(R_e+h)^2}=\frac{mv^2}{R_e+h}\\
\implies \frac{GM_e}{R_e+h}=v^2"

But, Radius of earth "R_e" is very larger than "h" , so we can apply the binomial approximation and thus

"v^2=\frac{GM_e}{R_e+h}=\frac{GM_e}{R_e(1+\frac{h}{R_e})}=\frac{GM_e}{R_e}(1+\frac{h}{R_e})^{-1}\\
\implies v^2\approx \frac{GM_e}{R_e}(1-\frac{h}{R_e})\approx g(R_e-h)\\
\implies v=\sqrt{g(R_e-h)}"

Thus, on plugin the value we get

"v=\sqrt{9.8(6400000-500000)}\approx7604m/s"


Since, for one complete revolution satellite covers

"d=2\pi R_e"

Thus, time takes to complete one revolution is

"T=\frac{d}{v}=\frac{2\pi R_e}{v}"

On plug int he data we get,

"T=5288.32s"